# Running in circles: practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

### v2

@@ 5 Acknowledgements
$X^{(1)}_{i}=X_{i}+\tau Z_{i}\qquad\text{and}\qquad X^{(2)}_{i}=X_{i}-\frac{1}{\tau}Z_{i}$
We can immediately deduce the marginal distributions of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ for all $i=1,\dots,n$, thanks to the independence between $X_{i}$ and $Z_{i}$:
@@
We can then derive the conditional distributions of $X^{(1)}_{i}|C_{1}$ and $X^{(1)}_{i}|C_{2}$. Indeed, $\mathbb{P}(X^{(1)}_{i}=x|C_{1})=\mathbb{P}(X^{(1)}_{i}=x|X^{(1)}_{i}>\mu)$ with $X^{(1)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}_{X^{(1)}}\right)$. Therefore, $X^{(1)}_{i}|X^{(1)}_{i}>\mu$ follows a half-normal distribution, and for all $i=1,\dots,n$:
@@
$X^{(2)}_{i}|X^{(1)}_{i}\sim\mathcal{N}\left(\mu+\frac{\rho\sigma_{X^{(1)}}\sigma_{X^{(2)}}}{\sigma^{2}_{X^{(1)}}}\left(X^{(1)}_{i}-\mu\right),\sigma^{2}_{X^{(2)}}-\frac{\rho^{2}\sigma^{2}_{X^{(1)}}\sigma^{2}_{X^{(2)}}}{\sigma^{2}_{X^{(1)}}}\right)$ (8)
@@
Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it does not contain any true clusters, and thus no real difference in means exists between subgroups of observations other than what is due to clustering. Therefore, the test statistic $T$ should be under $\mathcal{H}_{0}$, and thus centered around 00. In our result in (10), we observe a deviation of the distribution of $T$ from 00, quantified by $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$. The Type I error at level $\alpha$ associated with this test is then given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of the normal distribution $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the standard normal distribution $\mathcal{N}(0,1)$.
Here we assumed that all variances were known, and thus $s^{2}\left(X^{(2)}\right)$ was known as well. Therefore, it was possible to calculate the distribution of the test statistic for the $Z$ test to compare means between two samples. However, this result extends easily to the more practical case where $s^{2}\left(X^{(2)}\right)$ is unknown and an estimate $\widehat{s^{2}}\left(X^{(2)}\right)$ is used instead. In this case, still assuming that variances are equal between the two clusters, the distribution of the test statistic $T$ follows a Student’s $t$ distribution $\mathcal{T}(n-2)$ (due to the uncertainties associated with estimating this common variance). The associated Type I error then becomes: $1-F_{\mathcal{T}}(q_{\alpha/2})+F_{\mathcal{T}}(-q_{\alpha/2})$, where $F_{\mathcal{T}}$ is the cumulative distribution function of the non-central Student’s $t$ distribution with mean $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$ and $n-2$ degrees of freedom, and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the Student’s $t$ distribution with $n-2$ degrees of freedom.
This result underscores the crucial importance of precise variance estimation for the application of data fission. Indeed, for the test to be valid, that is, for the Type I error to be controlled at level $\alpha$, the distribution of the test statistic must be centered at 00. This implies that:

### v3

@@ 2.2 Marginal Data Fission
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}$
Appendix 4 Derivation of $t$-test statistic under the null of no-cluster in the univariate data fission post-clustering setting
Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Let $W_{1},\dots,W_{n}\overset{\text{iid}}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in(0,+\infty)$. Here, $b^{2}$ represents any value used as a plug-in for the variance of $X$, and in particular, $b^{2}$ can be an estimate obtained for $\sigma^{2}$. For all $i=1,\dots,n$, the splitting process of $X_{i}$ is given by:
$X^{(1)}_{i}=X_{i}+\tau W_{i}\qquad\text{and}\qquad X^{(2)}_{i}=X_{i}-\frac{1}{\tau}W_{i}$
We can immediately deduce the marginal distributions of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ for all $i=1,\dots,n$, thanks to the independence between $X_{i}$ and $W_{i}$:
@@
[…]
@@
[…]
@@
Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it does not contain any true clusters, and thus no real difference in means exists between subgroups of observations other than what is due to clustering. Therefore, the test statistic $T$ should be under $\mathcal{H}_{0}$, and thus centered around 0. In our result in (10), we observe a deviation of the distribution of $T$ from 0, quantified by $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$. The Type I error at level $\alpha$ associated with this test is then given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of the normal distribution $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the standard normal distribution $\mathcal{N}(0,1)$.
Here we assumed that all variances were known, and thus $s^{2}\left(X^{(2)}\right)$ was known as well. Therefore, it was possible to calculate the distribution of the test statistic for the $Z$ test to compare means between two samples. However, this result extends easily to the more practical case where $s^{2}\left(X^{(2)}\right)$ is unknown and an estimate $\widehat{s^{2}}\left(X^{(2)}\right)$ is used instead. In this case, still assuming that variances are equal between the two clusters, the distribution of the test statistic $T$ follows a Student’s $t$ distribution $\mathcal{T}(n-2)$ (due to the uncertainties associated with estimating this common variance). The associated Type I error then becomes: $1-F_{\mathcal{T}}(q_{\alpha/2})+F_{\mathcal{T}}(-q_{\alpha/2})$, where $F_{\mathcal{T}}$ is the cumulative distribution function of the non-central Student’s $t$ distribution with mean $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$ and n2n-2 degrees of freedom, and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the Student’s $t$ distribution with n2n-2 degrees of freedom.
This result underscores the crucial importance of precise variance estimation for the application of data fission. Indeed, for the test to be valid, that is, for the Type I error to be controlled at level $\alpha$, the distribution of the test statistic must be centered at 0. This implies that:
@@
Appendix 5 Supplementary Table 1
Conditional Fission Marginal Fission
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)$ $\sum\limits_{g=1}^{G}\pi_{g}\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)^{T}$ 0
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g\right)$ 0 $\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma}$
Table 1: Comparison of covariances under conditional and marginal data fission. The first row presents the covariance between the two fissioned random variables $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$, while the second row shows the conditional covariance given class membership $Z_{i}=g$. Conditional fission preserves within-component independence, whereas marginal fission introduces bias due to the discrepancy between component-specific and global covariance structures.
Appendix 6 Supplementary Figure 1
Figure 1: Schematic view illustrating the circularity induced by conditional data fission for post-clustering differential analysis.